Circular No. 6322

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVALIKE VARIABLE IN SAGITTARIUS S. Nakano, Sumoto, Japan, reports the discovery by Yukio Sakurai,

Otsuka-cho, Mito, Ibaragi-ken, of a possible “slow” nova, on Fuji G400 film taken with a 300-mm f/2.8 lens, located at α = 17^{h}52^{m}33^{s}.5, δ = −17^{o}40′52′′ (equinox 2000.0); the photograph taken on Feb. 20.806 UT shows the star as red in color and of mag 11.4. Patrol films taken by Sakurai during 1993–1994 show no candidate at this location, but the star is visible on films beginning in 1995 January (when it was of mag 12.5) and continuing through May, August–October, and again in 1996 January– February — the star slowly brightening over the past year. M. Hazen, Harvard-Smithsonian Center for Astrophysics, reports that a search of some 200 plates in the Harvard plate collection, rather evenly distributed in date of exposure during 1930–1951 and reaching blue mag 14 (for earlier plates) to 16 (later plates), shows no obvious variable star at this location. R. H. McNaught, Anglo-Australian Observatory, reports that a search of Schmidt films reveals “no strong indication of a red variable that can reach mag 12, but there are three red excess objects within ∼ 1′ of the quoted position. One is clearly variable but in the range R = 17–20 and is ∼ 1′ west of Sakurai’s position. Neither of the other red stars is obviously variable and are of (very roughly) R = 16 and 18.” He notes that a more precise position would be useful.

S. Benetti and H. W. Duerbeck, European Southern Observatory, report: “E. Cappellaro (Osservatorio Astronomico di Padova) observed this possible nova with the Dutch 0.9-m telescope at La Silla (ESO) on Feb. 23.3 UT, finding indeed a new star; preliminary photometry gives V = 12.8, B − V = +0.6, V − R = +0.6, V − I = +1.0. An inspection of a fullyreduced CCD spectrogram (range 375–985 nm; resolution 1.6 nm), taken by B. Leibundgut (ESO) with the ESO 3.6-m reflector (+ EFOSC1) on Feb. 23.4, reveals that the spectrum is consistent with a reddened early G-type star of high luminosity; no emission lines are visible. Hβ is in absorption at 485.7 nm, while Hα could be present in absorption, blended with another absorption feature (at 657.75 nm). Na I D lines are the strongest absorption feature of the spectrum (measured at 588.7 nm). The probable pre-outburst counterpart is visible as a star of m_{J} = 21, m_{R} = 20.5 on the ESO/SRC sky survey films. While the outburst amplitude and lightcurve suggest a slow or symbiotic nova, the lack of obvious emission lines one year after brightening is very unusual.”

1996 February 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green